# Mass Modeling and Kinematics of Galaxy Clusters in Modified Gravity

### v1

@@ 6 Discussions and conclusions
One can argue that the absence of biases due to the choice of the mass profile is just a consequence of the wide statistical uncertainties; as such, it is interesting to quantify the amount of systematic shift due to the wrong choice of mass model in an ideal situation where the statistical uncertainties are drastically reduced. A pps with more than $10^{4}$ particles within the virial radius following an NFW mass profile in GR was generated. A further perfect, unbiased, knowledge of the parameters $r_{200}$ and $r_{\text{s}}$ (i.e. an infinitely precise prior) [^11] One can think of this case as a sort of stacking of several clusters, even if in that case the effect of the mass bias would have been averaged out was assumed. The ($\phi_{2},\,\mathcal{Q}_{2}$) parameters are then constrained with the MG-MAMPOSSt code for the various mass density profile choices.
Also, in this case, one finds consistency with GR and no evident systematic bias due to the modelling. This indicates that the mass profile reconstruction is robust against the choice of the matter density profile. The main difference is, once again, the shape of the constraints related to the efficiency of the Chameleon screening. As an example, Fig. 7 shows the marginalized distribution of ($\phi_{2},`,\mathcal{Q}_{2}$) for the fit with an NFW model and a Burkert model (a similar trend is valid for the other profiles).
@@
From the posterior of Fig. 7 one gets $|f_{R,0}|\lesssim 4.4\times 10^{-6}$ for the NFW case and $|f_{R,0}|\lesssim 3.5\times 10^{-6}$ for the Burkert case, at $95\,\%$ C.L. upper limits. These bounds are in line with current cosmological constraints on the field (e.g. [81, 45, 50]) [^12] Note that tests at galactic scales [44] impose more stringent limits of $|f_{R,0}|\lesssim 10^{-8}$, when a particular model for the $f(R)$ function is adopted., and provide the limiting precision that can be achieved by kinematics and lensing analysis of clusters.
@@ Acknowledgements
BSH is supported by the INFN INDARK grant and acknowledges support from the COSMOS project of the Italian Space Agency (cosmosnet.it). CB acknowledges support from the COSMOS project of the Italian Space Agency (cosmosnet.it), and the INDARK Initiative of the INFN (web.infn.it /CSN4 /IS /Linea5/ InDark). A. M. Pombo is supported by the Czech Grant Agency (GAĈR) under grant number 21-16583M. AL has been supported by the EU H2020-MSCA-ITN-2019 Project 860744 ‘BiD4BESt: Big Data applications for Black Hole Evolution Studies’ and by the PRIN MIUR 2017 prot. 20173ML3WW, ‘Opening the ALMA window on the cosmic evolution of gas, stars, and supermassive black holes’.

### v2

@@ 6 Discussions and conclusions
One can argue that the absence of biases due to the choice of the mass profile is just a consequence of the wide statistical uncertainties; as such, it is interesting to quantify the amount of systematic shift due to the wrong choice of mass model in an ideal situation where the statistical uncertainties are drastically reduced. A pps with more than $10^{4}$ particles within the virial radius following an NFW mass profile in GR was generated. A further perfect, unbiased, knowledge of the parameters $r_{200}$ and $r_{\text{s}}$ (i.e. an infinitely precise prior, equivalent to set $\sigma_{r200},\sigma_{r_{\rm s}}\to 0$ in the lensing Gaussian distribution) [^11] One can think of this case as a sort of stacking of several clusters, even if in that case the effect of the mass bias would have been averaged out was assumed. The ($\phi_{2},\,\mathcal{Q}_{2}$) parameters are then constrained with the MG-MAMPOSSt code for the various mass density profile choices.
Also, in this case, one finds consistency with GR and no evident systematic bias due to the modelling. This indicates that the mass profile reconstruction is robust against the choice of the matter density profile. The main difference is, once again, the shape of the constraints related to the efficiency of the Chameleon screening. As an example, Fig. 9 shows the marginalized distribution of ($\phi_{2},\,\mathcal{Q}_{2}$) for the fit with an NFW model and a Burkert model (a similar trend is valid for the other profiles).
@@
Note that, although the results have been presented here for a single halo size, it has been checked that possible biases are negligible when changing the parameters $r_{200},r_{-2}$ (or equivalently $\rho_{\rm s},r_{\rm s}$ of the mock halo). However, the efficiency of the screening mechanism strongly depends also on the halo size, as shown in e.g. [66]; this affects the constrained region of ($\mathcal{Q}_{2},\phi_{2}$) in the joint kinematic+lensing analysis. In particular, low mass halos with smaller $r_{200}$ and with a lower concentration $c_{200}=r_{200}/r_{-2}$ (i.e. small values of the product $r_{\text{s}}^{2}\rho_{\text{s}}$) are characterized by overall less screening. This can be also visually spotted by looking at the screening equations in Sec. 3.
@@
From the posterior of Fig. 9 one gets $|f_{R,0}|\lesssim 4.4\times 10^{-6}$ for the NFW case and $|f_{R,0}|\lesssim 3.5\times 10^{-6}$ for the Burkert case, at $95\,\%$ C.L. upper limits. These bounds are in line with current cosmological constraints on the field (e.g. [85, 46, 51]) [^12] Note that tests at galactic scales [44] impose more stringent limits of $|f_{R,0}|\lesssim 10^{-8}$, when a particular model for the $f(R)$ function is adopted., and provide the limiting precision that can be achieved by kinematics and lensing analysis of clusters. It is worth to point out that the NFW model exhibits a bump in the allowed region at large $\mathcal{Q}_{2}$, which is a consequence of the stronger efficiency of the screening mechanism. Indeed, as one can visually inspect form Eq. 3.7, increasing the coupling $\beta$ results in an increase of the screening radius, suppressing the fifth force for reasonably small $\phi_{2}$ (see also the discussion in [66]).
@@ Acknowledgements
We thank the anonymous Referee for her/his precious comments and suggestions. BSH is supported by the INFN INDARK grant and acknowledges support from the COSMOS project of the Italian Space Agency (cosmosnet.it). CB acknowledges support from the COSMOS project of the Italian Space Agency (cosmosnet.it), and the INDARK Initiative of the INFN (web.infn.it /CSN4 /IS /Linea5/ InDark). A. M. Pombo is supported by the Czech Grant Agency (GAĈR) under grant number 21-16583M. AL has been supported by the EU H2020-MSCA-ITN-2019 Project 860744 ‘BiD4BESt: Big Data applications for Black Hole Evolution Studies’ and by the PRIN MIUR 2017 prot. 20173ML3WW, ‘Opening the ALMA window on the cosmic evolution of gas, stars, and supermassive black holes’.